1. Differentiate the following functions:

(a) y = x^3 cos


(b) y = e^x/sin 2x


1
Expert's answer
2021-05-11T13:58:15-0400

a). y=x3 cos x

"{dy\over dx }=f'g + fg'"

Let f = x3 "\implies" f'=3x2

Let g=cos(x) "\implies" g' =- sin(x)

"\therefore" "{dy\over dx}" = 3x2cos(x)-x3sin(x)

b). "y= {e^x\over sin(2x)}"

"{dy\over dx }=" "{f'g - fg'\over g^2}"

let "f = e^x" "\implies" "f' = e^x"

Let "g = sin(2x)" "\implies" "g' = 2 cos(2x)"

"\therefore" "{dy\over dx}=" "{e^x sin (2x) - e^x 2cos(2x)\over sin ^2 2x}"